A population has a standard deviation of 50. A random sample of 100 items from this population is selected. The sample mean is determined to be 600. At 95% confidence, the margin of error is
Answer by stanbon(57361)

A population has a standard deviation of 50.
A random sample of 100 items from this population is selected.
The sample mean is determined to be 600.
At 95% confidence, the margin of error is
-----------------------------------------------
E = ts/sqrt(n)
---
E = t(0.975, df= 99)*50/sqrt(100)
---
= 1.9842*50/10
---
= 9.9211..
=======================
Cheers,
Stan H.